A jeweler has 30 oz of an alloy consisting of 60 percent gold and 40 percent silver. How much of a second alloy containing 80 percent gold and 20 percent silver must be mixed with the first alloy to obtain an alloy containing 75 percent gold and 25 percent silver?
Answer by stanbon(75887)

A jeweler has 30 oz of an alloy consisting of 60 percent gold and 40 percent silver. How much of a second alloy containing 80 percent gold and 20 percent silver must be mixed with the first alloy to obtain an alloy containing 75 percent gold and 25 percent silver?
---------
Equation:
gold + gold = gold
-----
0.6*30 + 0.8x = 0.75*(30+x)
----------------------------------
60*30 + 80x = 75*30 + 75x
5x = 15*30
x = 90 oz (amt. of 80% needed)
-----------
Cheers,
Stan H.
--------------